Env. 6.6 The Department of Natural Resources (DNR) received a complaint from recreational fish- ermen that a community was releasing sewage into the river where they fished. These types of releases lower the level of dissolved oxygen in the river and hence cause damage to the fish resid ing in the river. An inspector from the DNR designs a study to investigate the fishermens claim. Fifteen water samples are selected at locations on the river upstream from the community and fifteen samples are selected downstream from the community. The dissolved oxygen readings in parts per million (ppm) are given in the following table Upstream Downstrea 3.2 3.4 3.7 3.9 3.63.8 3.9 3.6 .2 4.8 5.1 5.0 494.85.0 4.7 4.7 5.0 4.65.2 5.0 4.94.7 3.3 4.53.7 39 3.8 3.7
Write the 95% confidence interval for the true mean difference between the dissolved oxygen Upstream and Downstream (Hypstream-μDownstream). Interpret the confidence interval
Answer #1

Solution Given let x = Upstream X2 - Downstaeum (X-X) 2 (X2 72 72 3.2 5.2 4.8 5. 0.086 0.011 3.4 3.7 0.037 3.9 0.009 4.9 4.8ŽX2=561 X = 3.74 {(x2-x2)2 - 1.436 Sny El X-X2 E(X2=X2,2 n- = V 0.489 0.489 = V 1.436 14 14 Si=10.0349 0.0349 0.1025 5,= 0.18sp= 0.2621 trezni+N2-2) = t10.025.28) = 2:048 by table CI=(4.907-3.74 € (0.262))(2.098) A SIS = ( 1.167 + (0.2621) (2.048) (0